OpenStudy (anonymous):

If f(x) = 2x2 − 10, find f(5).

OpenStudy (anonymous):

place 5 in place of x

OpenStudy (hoblos):

put (5) in the place of x f(5) = 2(5)^2 -10 = 2(25) -10 =50-10 =40